OpenStudy (anonymous):

Given an isosceles triangle whose angle measures are 2x, 4x and 4x, calculate the value of x so that the sum of the angle measures is 180 degrees.

OpenStudy (anonymous):

You can make an equation: \[2x+4x+4x=180\] Then you can combine like terms... \[10x = 180\] Then ... \[\frac{ 10x }{ 10 }=\frac{ 180 }{ 10 }=18 = x\] The value of x = 18. And you can check this by substituting x with 18 ... \[2(18)+4(18)+4(18)=180\] Then once you simplify the above, you get ... 180 = 180 So, the value of x = 18. That's how I do it.